HIGH RESOLUTION, NEAR AND MID IR, JET-COOLED, CAVITY RING-DOWN SPECTROSCOPY WITH A NOVEL DISCHARGE EXPANSION

\underline{SHENGHAI WU}, PATRICK DUPRÉ AND TERRY A MILLER, Department of Chemistry, The Ohio State University, 120 W. 18^{th} Ave., Columbus OH, 43210.

The frequency coverage of our CRDS spectrometer^{a} has been extended from the 1.0 – 1.5 μm region to the 3 – 3.5 μm region, which opens exciting opportunities for studying the C-H fundamental vibrational modes of molecules and free radicals. Furthermore, higher resolution has been provided by using difference frequency generation (DFG) based on a Fourier transform limited Ti:Sapphire pulse-amplified laser^{b}. As an example, the jet-cooled spectra of the CH_{3} radical with partially resolved hyperfine structures will be presented. Non-exponential effects in the ring-down decay have also been minimized by using the narrow bandwidth radiation. Additionally, a novel transverse discharge arrangement had been designed to improve stability, and also to provide a capability of mixing chemical reagents downstream. We will present preliminary spectral results and kinetic analyses characterizing this radical source.

aShenghai Wu, Patrick Dupré and Terry A Miller, Phys. Chem. Chem. Phys., 2006 (Accepted, DOI: 10.1039/b518279d)

bPatrick Dupré and Terry A Miller, The 60th Symposium on Molecular Spectroscopy, 2005, TJ10